# Recent advances in text embedding: A Comprehensive Review of Top-Performing Methods on the MTEB Benchmark

### v1

@@ Abstract
Text embedding methods have become increasingly popular in both industrial and academic fields due to their critical role in a variety of natural language processing tasks. The significance of universal text embeddings has been further highlighted with the rise of Large Language Models (LLMs) applications such as Retrieval-Augmented Systems (RAGs). While previous models have attempted to be general-purpose, they often struggle to generalize across tasks and domains. However, recent advancements in training data quantity, quality and diversity; synthetic data generation from LLMs as well as using LLMs as backbones encourage great improvements in pursuing universal text embeddings. In this paper, we provide an overview of the recent advances in universal text embedding models with a focus on the top performing text embeddings on Massive Text Embedding Benchmark (MTEB). Through detailed comparison and analysis, we highlight the key contributions and limitations in this area, and propose potentially inspiring future research directions.
@@ 1 Introduction
Text embedding methods have gained considerable interest in both industry and academia due to their important role in various natural language processing tasks such as text classification (?), text clustering (?, ?), sentiment analysis (?, ?), information retrieval (?), question answering (?), dialogue systems (?), semantic textual similarity (?), item recommendation (?) and so on (?, ?, ?). With the increasing popularity of Large Language Models (LLMs) based applications such as Retrieval-Augmented Systems (RAGs), the pivotal role of text embeddings has been underscored recently. This is mainly due to the fact that these LLM based applications are heavily dependent on the high quality text embeddings for tasks like vector search, a process where the most relevant documents are retrieved for LLM Question Answering (QA) (?, ?). Source attribution of generated text is another important application of text embeddings (?) that can improve the interpretability and trustworthiness of LLMs (?).
@@
  • 1st era: Count-based Embeddings. Bag of Words and Term Frequency-Inverse Document Frequency (TF-IDF) are two representative works in this era. Bag of Words (?) is one of the earliest text representation methods, which counts the presence or occurrence of each word in the documents and use them as features. TF-IDF measures how important a word is to a document relative to a corpus, by increasing proportionally to the number of times a word appears in the document but offset by the frequency of the word in the corpus (?). Both BoW and TF-IDF highlight the word/term relevancy instead of using the context information or the meaning of words (?). There are also other works in this era transforming texts into low-dimensional dense embeddings such as Latent Semantic Indexing (LSA) (?) (generating document embeddings with the decomposition of a word-document co-occurrence matrix) (?).
@@
  • 2nd era: Static dense word embeddings. Word2Vec (?), GloVe (?) and FastText (?) are representative works that showed a significant step forward in the field of text representations using the surroundings of words to generate dense vector representations. Word2Vec focuses on local context using either Continuous Bag of Words (CBOW) approach (given the context, it predicts the target word) or Skip-gram approach (given the word, it predicts the context). Instead of only focusing on local context like Word2Vec, GloVe also takes the global corpus statistics into account. FastText further improves word embeddings by capturing the internal structure or morphology of words with a focus on character-level information of words and learning representations of sub-words (?). Even though these models can capture a range of semantic and syntactic similarities successfully, they provide a single static vector per word, which ignores the fact that a word’s meaning can be influenced by its surrounding context.
@@
  • 3rd era: Contextualized embeddings. The third era of text embeddings ushers in a new phase of embedding sophistication: context-sensitive dynamic embeddings that adapt or change based on context. Representative works include Embeddings From Language Models (ELMo) (?), Generative Pre-trained Transformer (GPT) (?) and Bidirectional Encoder Representations from Transformers (BERT) (?). ELMo models the polysemy using a bidirectional Long Short Term Memory Network (LSTM) with the concatenation of the left-to-right and right-to-left representations. Unlike ELMo, GPT uses Transformer (one-way instead of bi-directional) (?) to learn the text embedding using a combination of unsupervised pre-training and supervised fine-tuning. It was observed that attentional memory of the transformer assisted in (better) transfer compared to LSTMs (?). BERT instead uses a bidirectional Transformer encoder to take into account both the left and right context for Masked Language Model (MLM) and Next Sentence Prediction (NSP) tasks during pre-training, which allows for a deeper understanding of word relationships by considering the full context of a word in a sentence in both directions. (?, ?).
@@
  • 4th era: Universal text embeddings. The pursuit of developing a unified model to address a multitude of downstream tasks has been long-standing (?). Despite attempting to be general-purpose in previous models such as (?, ?, ?), studies indicate that these embedding models struggle to generalize across tasks and domains (?). Thanks to the increasing number and improved quality of diverse text datasets across different tasks (?, ?), good quality synthetic data generated by LLMs (?, ?) as well as benchmarks with the focus on novel task and domain generalization such as the Massive Text Embedding Benchmark (MTEB) (?); the universality of text embeddings can be improved and evaluated across various languages and tasks such as retrieval, ranking, clustering, among others. The creation of unified models trained across diverse tasks has started to make progress with representative works like GTE (?), BGE (?), E5 (?, ?, ?), Gecko (?), LLM2Vec (?), etc.
@@
There are several reviews on text embeddings such as in (?, ?, ?, ?, ?), but none of the existing work focus on the recent advances in the universal text embeddings in the fourth era. To fill in the gap, the main focus of this work is to review recent advances in universal text embeddings. More specifically, the top performing text embeddings in the Massive Text Embedding Benchmark (MTEB) (?) are the main focus of this review. The remainder of this paper is organized as follows: the preliminaries, background and categorization of 4th era universal text embeddings are introduced in Section 2. In Section 3, 4 and 5, the overview of the top performing state of the art text embeddings and their main contributions are explained. We describe the trends, performance and efficiency analysis of the state of the art text embeddings as well as their limitations in Section 6. Finally, the conclusion and future directions in text embeddings are given in Section 7.
@@ 2.1 Definitions
In the context of Natural Language Processing (NLP) or Natural Language Understanding (NLU), text refers to a collection of words, phrases, sentences, paragraphs or larger utterance that convey meaningful information (?). The form and length of text often vary on the task such as text classification/clustering, sentiment analysis, information retrieval, dialogue systems, item recommendation, etc. However, an embedding is a fixed-length low-dimensional dense vector representation (?). Text embedding then can be defined as a numerical dense representation of a word, phrase, sentence, or larger utterance in natural language in a certain space where texts with similar meanings are near each other (?, ?, ?, ?). The meaning of a word is influenced by its context, and it is from this context that a word embedding is usually learnt. The meaning of a sentence is more complex because it depends on the words used in the sentence, the syntactic structure as well as the surrounding sentences (?). The meaning of a document is even more complex as it is a high-level abstraction of the whole text (words, sentences, paragraphs, etc.). The definition of ”meaning”, ”local information” or ”context” changes when the text length changes, which makes it a great challenge to learn the embedding for an ”arbitrary span of contiguous text” (?).
@@
In recent works, universal text embedding (?, ?, ?) or general-purpose text embedding as used in (?, ?, ?, ?) generally means a unified comprehensive text embedding model that can address a multitude of downstream tasks. In other words, the universal text embedding is not just proficient in a single particular task, but it proves to be consistently beneficial across a range of tasks such as text classification, text clustering, sentiment analysis, semantic textual similarity, summarization, retrieval tasks, etc. The objective of creating universal text embeddings is to mimic the fundamental process of how humans understand and process text, which can be beneficial in various domains (?). With the recent work such as (?, ?), the definition of universal text embedding has been extended to multi-task, multi-lingual, while (?) shows that a natural language model can also understand well programming languages. In this work, we define universal text embedding as a unified comprehensive text embedding model that can address a multitude of input text length, downstream tasks, domains and languages. The research of universal text embedding has been stimulated by several recent developments. These include the growth in quantity and refinement in quality of diverse text datasets across various tasks (?, ?), the production of high-quality synthetic data by LLMs (?, ?), and benchmarks that emphasize new task and domain generalization, such as the multi-lingual Massive Text Embedding Benchmark (MTEB) (?).
@@ 2.2 Background
In this work, we study and analyze the top performing text embedding models that are either open-source or well documented from MTEB English benchmark (because the English benchmark has more and diverse evaluation tasks compared to other languages). It can be found that BERT-based models used in (?, ?, ?, ?, ?) and LLMs used in (?, ?, ?, ?, ?, ?) are two most popular backbones of the top performing universal text embedding models on the MTEB English benchmark.
BERT (?)
To generate contextual embeddings, BERT, pre-trained on a massive corpus and fine-tuned using labeled data from the downstream tasks, employs a bidirectional Transformer encoder to take into account both the left and right context in all layers. To alleviate the uni-directionality constraint, BERT proposes a masked language modelling (MLM) objective, where some of the tokens of a input sequence are randomly masked, and the objective is to predict the vocab-ids of the masked tokens based only on its context (?). Additionally, a Next Sentence Prediction (NSP) task is also used to jointly pre-train text-pair representations with the objective to help tasks that require reasoning over text pairs (?). WordPiece embeddings with a 30,000 tokens vocabulary (?) is used by BERT, with special tokens including [CLS] token (a special classification token as the first token of each sequence) and [SEP] token to separate sentence pairs. The final hidden state of [CLS] is used for sentence-level tasks and the final hidden state of each token is used for token-level tasks (?, ?). Some important details about BERT include:
  • Pre-training data: BooksCorpus (800M words) (?) and English Wikipedia ignoring lists, tables, and headers (2,500M words).
@@
  • Loss function: the sum of the mean MLM likelihood and the mean NSP likelihood (?).
@@
Following the success of BERT, several BERT-based models have been introduced, such as Robustly Optimized BERT Pretraining Approach (RoBERTa) (?), Distilled version of BERT (DistilBERT) (?), and A Lite BERT (ALBERT) (?), each offering unique enhancements and optimizations while maintaining the core bidirectional approach of the original BERT model. One of the limitations of the BERT network structure is that no independent sentence embeddings are computed, which makes it difficult to use for various pair regression tasks due to large number of combinations. To allow for more efficient sentence-level embeddings, Sentence-BERT (SBERT) introduces the siamese and triplet network structures to generate highly effective semantically meaningful sentence embeddings that can be compared with cosine similarity, which has served as a cornerstone for further research (?, ?). Another cornerstone work is Simple Contrastive Learning of Sentence Embeddings (SimCSE) (?) using unsupervised and supervised contrastive learning, which is widely adopted by recent state of the art text embeddings.
@@
The widespread use of ChatGPT has showcased the impressive abilities of Large Language Models (LLMs) in following instructions, in-context learning with minimal few-shot examples and amazing conversation abilities with humans. While some of the best performing LLMs like GPT-4 (?) are proprietary with limited technical information available, some open-source LLM models like LLaMA-2 (?), LLaMA-3 (?) and Mistral (?) have made some notable efforts to catch up(?). One advantage of using LLMs for text embedding is that they are extensively pre-trained on web-scale data already, which does not need the contrastive pre-training step used in existing state of the art text embedding models. At present, the foundation for the majority of LLMs is the Transformer architecture, which employs layers of multi-head attention in a very deep neural network. Decoder-only LLMs utilize the causal attention mechanism, where the representation of a token at a specific position $i$ is exclusively impacted by the representations of tokens that come before it. The authors from (?) hypothesize that causal attention mechanism might partly be the reason of the slow adoption of decoder-only LLMs for text embedding tasks as it inherently limits their ability to produce rich contextualized representations. Several recent works such (?, ?, ?, ?, ?, ?) have proposed several solutions to mitigate such limitations.
@@
The objective of MTEB is to provide comprehensive understandings on the universality of text embedding models, including 58 datasets covering 112 languages from 8 embedding tasks: Bitext mining, Classification, Clustering, Pair classification, Reranking, Retrieval, Semantic Textual Similarity (STS) and Summarization (?). The leader-board results are available on the Hugging Face Hub [^1] https://huggingface.co/spaces/mteb/leaderboard, where the results of English (56 datasets), Chinese (35 datasets), French (26 datasets) and Polish (26 datasets) benchmark results can be found respectively.
@@ 2.3 Taxonomy of universal text embeddings
In this section, the main focuses and contributions of the some of the MTEB top performing state of the art text embedding methods are analyzed (shown in Figure 2), including: E5: EmbEddings from bidirEctional Encoder rEpresentations (?), GTE: General-purpose Text Embedding model (?), BGE: Beijing Academy of Artificial Intelligence (BAAI) General Embedding (?), UAE: Universal AnglE Embedding (?), MRL: Matryoshka Representation Learning (?), 2DMSE: 2D Matryoshka Sentence Embeddings (?), GRIT: Generative Representational Instruction Tuning (?), LLM2Vec: (?), Multilingual E5: (?), E5-mistral-7b-instruct: (?), Gecko: (?), Echo-mistral: (?), SFR-Embedding-Mistral: (?). The main focus/contributions are summarized and simplified as the following 4 aspects:
  • Real world data: one way to learn the universal text embedding is using a multi-stage contrastive learning strategy with diverse training data mixture. For example, GTE (?) uses diverse datasets for both pre-training and fine-tuning stage. BGE (?) introduces a compressive data package C-Pack, while E5 (?) constructed a curated web-scale text pair dataset named Colossal Clean text Pairs (CCPairs) containing heterogeneous training signals by combining various semi-structured data sources along with aggressive filtering (270M text pairs filtered from 1.3B noisy text pairs) with a consistency-based filter to improve data quality (?). Some works like GISTEmbed (?) also focus on improving the quality of hard negatives used for training.
  • Loss function: another research direction is to focus on improving the loss functions. As many existing text embedding works employed the cosine function in their training objective to measure the pairwise semantic similarity, the authors from UAE (?) point out that there is the gradient vanishing issue due to the saturation zones of cosine function, which hinder the ability to learn subtle distinctions between texts in back propagation. Hence they propose a novel angle-optimized text embedding model called AnglE with angle optimization in a complex space which substantially improve the text embedding quality in various scenarios. Matryoshka Representation Learning (MRL) (?) and 2D Matryoshka Sentence Embeddings (2DMSE) propose new loss functions in order to reduce the computational cost of downstream tasks.
@@
    • 1. use synthetic data generated by LLMs: In (?), the authors apply LLMs as data annotators to label the pseudo-supervised data for the training to improve the model performance. (?) and (?) use proprietary LLMs including GPT-35-Turbo and GPT-4 to generate synthetic data covering a various range of text embedding tasks in 93 languages (among which 25% are generated by GPT-35-Turbo and others are generated by GPT-4) to increase the training data diversity, while (?) use synthetic data generation to distill knowledge from large language models into a retriever.
    • 2. use LLMs as backbone for text embeddings: as LLMs are extensively pre-trained on web-scale data already, which does not need the large scale contrastive pre-training step used in existing state of the art text embedding models, many works also try to get embeddings directly from LLMs. For example, E5-mistral-7b-instruct perform multi-task fine-tuning on Mistral 7b model which is one of the best performing method on MTEB. Echo-mistral (?), LLM2Vec (?) and GRIT (?) propose various different ways so that decoder only LLMs can generate high quality text embeddings using bidirectional attention.

### v2

# Recent advances in universal text embeddings: A Comprehensive Review of Top-Performing Methods on the MTEB Benchmark

@@ Abstract
Text embedding methods have become increasingly popular in both industrial and academic fields due to their critical role in a variety of natural language processing tasks. The significance of universal text embeddings has been further highlighted with the rise of Large Language Models (LLMs) applications such as Retrieval-Augmented Systems (RAGs). While previous models have attempted to be general-purpose, they often struggle to generalize across tasks and domains. However, recent advancements in training data quantity, quality and diversity; synthetic data generation from LLMs as well as using LLMs as backbones encourage great improvements in pursuing universal text embeddings. In this paper, we provide an overview of the recent advances in universal text embedding models with a focus on the top performing text embeddings on Massive Text Embedding Benchmark (MTEB). Through detailed comparison and analysis, we offer a systematic organization of the literature, underscoring the significant developments and limitations in the recent advancements of universal text embedding models, and suggest potential future research directions that could inspire further advancements in this field.
{IEEEkeywords}
Language Models, Representation learning, Text embedding, Universal text embeddings
@@ 1 Introduction
Text embedding methods have gained considerable interest in both industry and academia due to their important role in various natural language processing tasks such as text classification [1], text clustering [2, 3], sentiment analysis [4, 5], information retrieval [6], question answering [7], dialogue systems [8], semantic textual similarity [9], item recommendation [10] and so on [11, 12, 13]. With the increasing popularity of Large Language Models (LLMs) based applications such as Retrieval-Augmented Systems (RAGs), the pivotal role of text embeddings has been underscored recently. This is mainly due to the fact that these LLM based applications are heavily dependent on the high quality text embeddings for tasks like vector search, a process where the most relevant documents are retrieved for LLM Question Answering (QA) [12, 14]. Source attribution of generated text is another important application of text embeddings [15] that can improve the interpretability and trustworthiness of LLMs [16].
@@
  • 1st era: Count-based Embeddings. Bag of Words and Term Frequency-Inverse Document Frequency (TF-IDF) are two representative works in this era. Bag of Words [17] is one of the earliest text representation methods, which counts the presence or occurrence of each word in the documents and use them as features. TF-IDF measures how important a word is to a document relative to a corpus, by increasing proportionally to the number of times a word appears in the document but offset by the frequency of the word in the corpus [18]. Both BoW and TF-IDF highlight the word/term relevancy instead of using the context information or the meaning of words [19]. There are also other works in this era transforming texts into low-dimensional dense embeddings such as Latent Semantic Indexing (LSA) [20] generating document embeddings with the decomposition of a word-document co-occurrence matrix [21].
@@
  • 2nd era: Static dense word embeddings. Word2Vec [22], GloVe [23] and FastText [24] are representative works that showed a significant step forward in the field of text representations using the surroundings of words to generate dense vector representations. Word2Vec focuses on local context using either Continuous Bag of Words (CBOW) approach (given the context, it predicts the target word) or Skip-gram approach (given the word, it predicts the context). Instead of only focusing on local context like Word2Vec, GloVe also takes the global corpus statistics into account. FastText further improves word embeddings by capturing the internal structure or morphology of words with a focus on character-level information of words and learning representations of sub-words [25]. Even though these models can capture a range of semantic and syntactic similarities successfully, they provide a single static vector per word, which ignores the fact that a word’s meaning can be influenced by its surrounding context.
@@
  • 3rd era: Contextualized embeddings. The third era of text embeddings ushers in a new phase of embedding sophistication: context-sensitive dynamic embeddings that adapt or change based on context. Representative works include Embeddings From Language Models (ELMo) [26], Generative Pre-trained Transformer (GPT) [27] and Bidirectional Encoder Representations from Transformers (BERT) [28]. ELMo models the polysemy using a bidirectional Long Short Term Memory Network (LSTM) with the concatenation of the left-to-right and right-to-left representations. Unlike ELMo, GPT uses Transformer (one-way instead of bi-directional) [29] to learn the text embedding using a combination of unsupervised pre-training and supervised fine-tuning. It was observed that attentional memory of the transformer assisted in (better) transfer compared to LSTMs [25]. BERT instead uses a bidirectional Transformer encoder to take into account both the left and right context for Masked Language Model (MLM) and Next Sentence Prediction (NSP) tasks during pre-training, which allows for a deeper understanding of word relationships by considering the full context of a word in a sentence in both directions. [28, 30].
@@
  • 4th era: Universal text embeddings. The pursuit of developing a unified model to address a multitude of downstream tasks has been long-standing [12]. Despite attempting to be general-purpose in previous models such as [31, 32, 33], studies indicate that these embedding models struggle to generalize across tasks and domains [34]. Thanks to the increasing number and improved quality of diverse text datasets across different tasks [35, 36], good quality synthetic data generated by LLMs [34, 16] as well as benchmarks with the focus on novel task and domain generalization such as the Massive Text Embedding Benchmark (MTEB) [37]; the universality of text embeddings can be improved and evaluated across various languages and tasks such as retrieval, ranking, clustering, among others. The creation of unified models trained across diverse tasks has started to make progress with representative works like GTE [12], BGE [35], E5 [21, 16, 38], Gecko [34], LLM2Vec [39], etc.
@@
There are several reviews on text embeddings such as in [40, 41, 25, 42, 43, 44], but none of the existing work focus on the recent advances in the universal text embeddings in the fourth era. To fill in the gap, the main focus of this work is to review recent advances in universal text embeddings. More specifically, the top performing text embeddings in the Massive Text Embedding Benchmark (MTEB) [37] are the main focus of this work. This survey offers a systematic organization of the literature, underscoring the significant developments and challenges in the recent advancements of universal text embedding models (primarily focusing on the methods proposed in years 2023 and 2024). Furthermore, we suggest potential future research directions that could inspire further advancements in this field. The reminder of this paper is organized as follows: the preliminaries, background and categorization of 4th era universal text embeddings are introduced in Section 2. In Section 3, 4 and 5, the overview of the top performing state of the art text embeddings and their main contributions are explained. We describe the trends, performance and efficiency analysis of the state of the art text embeddings as well as their limitations in Section 6. Finally, the conclusion and future directions in text embeddings are given in Section 7.
@@ 2.1 Definitions
In the context of Natural Language Processing (NLP) or Natural Language Understanding (NLU), text refers to a collection of words, phrases, sentences, paragraphs or larger utterance that convey meaningful information [45]. The form and length of text often vary on the task such as text classification/clustering, sentiment analysis, information retrieval, dialogue systems, item recommendation, etc. However, an embedding is a fixed-length low-dimensional dense vector representation [21]. Text embedding then can be defined as a numerical dense representation of a word, phrase, sentence, or larger utterance in natural language in a certain space where texts with similar meanings are near each other [3, 34, 43, 12]. The meaning of a word is influenced by its context, and it is from this context that a word embedding is usually learnt. The meaning of a sentence is more complex because it depends on the words used in the sentence, the syntactic structure as well as the surrounding sentences [46]. The meaning of a document is even more complex as it is a high-level abstraction of the whole text (words, sentences, paragraphs, etc.). The definition of ”meaning”, ”local information” or ”context” changes when the text length changes, which makes it a great challenge to learn the embedding for an ”arbitrary span of contiguous text” [28].
@@
In recent works, universal text embedding [35, 39, 47] or general-purpose text embedding as used in [34, 12, 21, 48] generally means a unified comprehensive text embedding model that can address a multitude of downstream tasks. In other words, the universal text embedding is not just proficient in a single particular task, but it proves to be consistently beneficial across a range of tasks such as text classification, text clustering, sentiment analysis, semantic textual similarity, summarization, retrieval tasks, etc. The objective of creating universal text embeddings is to mimic the fundamental process of how humans understand and process text, which can be beneficial in various domains [46]. With the recent work such as [38, 49], the definition of universal text embedding has been extended to multi-task, multi-lingual, while [12] shows that a natural language model can also understand well programming languages. In this work, we define universal text embedding as a unified comprehensive text embedding model that can address a multitude of input text length, downstream tasks, domains and languages. The research of universal text embedding has been stimulated by several recent developments. These include the growth in quantity and refinement in quality of diverse text datasets across various tasks [35, 36], the production of high-quality synthetic data by LLMs [34, 16], and benchmarks that emphasize new task and domain generalization, such as the multi-lingual Massive Text Embedding Benchmark (MTEB) [37].
@@ 2.2 Background
In this work, we study and analyze the top performing text embedding models that are either open-source or well documented from MTEB English benchmark (because the English benchmark has more and diverse evaluation tasks compared to other languages). It can be found that BERT-based models used in [12, 35, 21, 38, 50] and LLMs used in [38, 49, 39, 51, 48, 34] are two most popular backbones of the top performing universal text embedding models on the MTEB English benchmark.
BERT [28]
To generate contextual embeddings, BERT, pre-trained on a massive corpus and fine-tuned using labeled data from the downstream tasks, employs a bidirectional Transformer encoder to take into account both the left and right context in all layers. To alleviate the uni-directionality constraint, BERT proposes a masked language modelling (MLM) objective, where some of the tokens of a input sequence are randomly masked, and the objective is to predict the vocab-ids of the masked tokens based only on its context [28]. Additionally, a Next Sentence Prediction (NSP) task is also used to jointly pre-train text-pair representations with the objective to help tasks that require reasoning over text pairs [43]. WordPiece embeddings with a 30,000 tokens vocabulary [52] is used by BERT, with special tokens including [CLS] token (a special classification token as the first token of each sequence) and [SEP] token to separate sentence pairs. The final hidden state of [CLS] is used for sentence-level tasks and the final hidden state of each token is used for token-level tasks [28, 43]. Some important details about BERT include:
  • Pre-training data: BooksCorpus (800M words) [53] and English Wikipedia ignoring lists, tables, and headers (2,500M words).
@@
  • Loss function: the sum of the mean MLM likelihood and the mean NSP likelihood [28].
@@
Following the success of BERT, several BERT-based models have been introduced, such as Robustly Optimized BERT Pretraining Approach (RoBERTa) [54], Distilled version of BERT (DistilBERT) [55], and A Lite BERT (ALBERT) [56], each offering unique enhancements and optimizations while maintaining the core bidirectional approach of the original BERT model. One of the limitations of the BERT network structure is that no independent sentence embeddings are computed, which makes it difficult to use for various pair regression tasks due to large number of combinations. To allow for more efficient sentence-level embeddings, Sentence-BERT (SBERT) introduces the siamese and triplet network structures to generate highly effective semantically meaningful sentence embeddings that can be compared with cosine similarity, which has served as a cornerstone for further research [3, 44]. Another cornerstone work is Simple Contrastive Learning of Sentence Embeddings (SimCSE) [57] using unsupervised and supervised contrastive learning, which is widely adopted by recent state of the art text embeddings.
@@
The widespread use of ChatGPT has showcased the impressive abilities of Large Language Models (LLMs) in following instructions, in-context learning with minimal few-shot examples and amazing conversation abilities with humans. While some of the best performing LLMs like GPT-4 [58] are proprietary with limited technical information available, some open-source LLM models like LLaMA-2 [59], LLaMA-3 [60] and Mistral [61] have made some notable efforts to catch up[16]. One advantage of using LLMs for text embedding is that they are extensively pre-trained on web-scale data already, which does not need the contrastive pre-training step used in existing state of the art text embedding models. At present, the foundation for the majority of LLMs is the Transformer architecture, which employs layers of multi-head attention in a very deep neural network. Decoder-only LLMs utilize the causal attention mechanism, where the representation of a token at a specific position $i$ is exclusively impacted by the representations of tokens that come before it. The authors from [39] hypothesize that causal attention mechanism might partly be the reason of the slow adoption of decoder-only LLMs for text embedding tasks as it inherently limits their ability to produce rich contextualized representations. Several recent works such [38, 49, 39, 51, 48, 34] have proposed several solutions to mitigate such limitations.
@@
The objective of MTEB is to provide comprehensive understandings on the universality of text embedding models, including 58 datasets covering 112 languages from 8 embedding tasks: Bitext mining, Classification, Clustering, Pair classification, Reranking, Retrieval, Semantic Textual Similarity (STS) and Summarization [37]. The leader-board results are available on the Hugging Face Hub [^1] https://huggingface.co/spaces/mteb/leaderboard, where the results of English (56 datasets), Chinese (35 datasets), French (26 datasets) and Polish (26 datasets) benchmark results can be found respectively.
@@ 2.3 Taxonomy of universal text embeddings
In this section, the main focuses and contributions of the some of the MTEB top performing state of the art text embedding methods are analyzed (shown in Figure 2), including: E5: EmbEddings from bidirEctional Encoder rEpresentations [21], GTE: General-purpose Text Embedding model [12], BGE: Beijing Academy of Artificial Intelligence (BAAI) General Embedding [35], UAE: Universal AnglE Embedding [50], MRL: Matryoshka Representation Learning [62], 2DMSE: 2D Matryoshka Sentence Embeddings [63], GRIT: Generative Representational Instruction Tuning [49], LLM2Vec: [39], Multilingual E5: [38], E5-mistral-7b-instruct: [16], Gecko: [34], Echo-mistral: [48], SFR-Embedding-Mistral: [51]. The main focus/contributions are summarized and simplified as the following 4 aspects:
  • Real world data: one way to learn the universal text embedding is using a multi-stage contrastive learning strategy with diverse training data mixture. For example, GTE [12] uses diverse datasets for both pre-training and fine-tuning stage. BGE [35] introduces a compressive data package C-Pack, while E5 [21] constructed a curated web-scale text pair dataset named Colossal Clean text Pairs (CCPairs) containing heterogeneous training signals by combining various semi-structured data sources along with aggressive filtering (270M text pairs filtered from 1.3B noisy text pairs) with a consistency-based filter to improve data quality [64]. Some works like GISTEmbed [47] also focus on improving the quality of hard negatives used for training.
  • Loss function: another research direction is to focus on improving the loss functions. As many existing text embedding works employed the cosine function in their training objective to measure the pairwise semantic similarity, the authors from UAE [50] point out that there is the gradient vanishing issue due to the saturation zones of cosine function, which hinder the ability to learn subtle distinctions between texts in back propagation. Hence they propose a novel angle-optimized text embedding model called AnglE with angle optimization in a complex space which substantially improve the text embedding quality in various scenarios. Matryoshka Representation Learning (MRL) [62] and 2D Matryoshka Sentence Embeddings (2DMSE) propose new loss functions in order to reduce the computational cost of downstream tasks.
@@
    • 1. use synthetic data generated by LLMs: In [50], the authors apply LLMs as data annotators to label the pseudo-supervised data for the training to improve the model performance. [16] and [38] use proprietary LLMs including GPT-35-Turbo and GPT-4 to generate synthetic data covering a various range of text embedding tasks in 93 languages (among which 25% are generated by GPT-35-Turbo and others are generated by GPT-4) to increase the training data diversity, while [34] use synthetic data generation to distill knowledge from large language models into a retriever.
    • 2. use LLMs as backbone for text embeddings: as LLMs are extensively pre-trained on web-scale data already, which does not need the large scale contrastive pre-training step used in existing state of the art text embedding models, many works also try to get embeddings directly from LLMs. For example, E5-mistral-7b-instruct perform multi-task fine-tuning on Mistral 7b model which is one of the best performing method on MTEB. Echo-mistral [48], LLM2Vec [39] and GRIT [49] propose various different ways so that decoder only LLMs can generate high quality text embeddings using bidirectional attention.